OpenStudy (anonymous):

find the maximum y value on the graph y=f(x) f(y)=-x^2+8x+6

OpenStudy (mathstudent55):

What math are you in? Algebra?

OpenStudy (mathstudent55):

An equation of this type is of a parabola that opens downward

OpenStudy (anonymous):

algebra 2 from what I understand

OpenStudy (mathstudent55):



OpenStudy (mathstudent55):

Since this is parabola that opens downward, you're looking for the vertex which is on top. That's the maximum y

OpenStudy (anonymous):

sorry my system went out so since the x^2 and 8x are x does that mean that the high point for y is 6

OpenStudy (mathstudent55):

For parabola ax^2 + bx + c = 0, -b/(2a) gives the x-coordinate of the vertex.

OpenStudy (anonymous):

so ot would be -1x^2+8cx+6=0 but then I do not lknow what to do

OpenStudy (mathstudent55):